Circular No. 6868

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0421+560 M. Orlandini, D. Dal Fiume, and F. Frontera, Istituto Tecnologie e Stu-

dio Radiazioni Extraterrestri, CNR, Bologna; L. A. Antonelli, BeppoSAX Science Data Center, Rome; L. Piro, Istituto Astrofisica Spaziale, CNR, Rome; and A. N. Parmar, Space Science Department, ESTEC, report: “On Apr. 3.2 UT, BeppoSAX observed XTE J0421+560 with its Narrow Field Instruments. The source position as determined by LECS and MECS is α = 4^{h}19^{m}46^{s}.0, δ = +55^{o}59′24′′ (equinox 2000.0; error radius 50′′), consistent with the Marshall et al. position (IAUC 6857) and the CI Cam position (IAUC 6862). The source was fading during the 45 000-s observation; a fit with an exponential decay law gives an e-folding time of 0.89 ± 0.02 day. We observe a prominent iron line in emission at 6.7 keV with an equivalent width of 670 eV, as well as other low-energy features. The 2–10- and 15– 100-keV fluxes averaged over the whole observation are 1.16 × 10^{−9} and 1.58 × 10^{−10} erg cm^{−2} s^{−1}, respectively.”

GRB 980329 I. A. Smith, Rice University; and R. P. J. Tilanus, Joint Astronomy

Centre, report on behalf of the James Clerk Maxwell Telescope (JCMT) GRB collaboration: “We used the JCMT (+ SCUBA submillimeter continuum bolometer array) to observe the variable radio source VLA J0702+3850 proposed by Taylor et al. (see IAUC 6864) as the radio afterglow to GRB 980329. On Apr. 5.2 UT, we detected a source at 850 μm with a preliminary flux density of 5 ± 1.5 mJy. This source was confirmed on Apr. 6.2 with a flux density of 4 ± 1.2 mJy, resulting in an average of 4.5 ± 1 mJy over the two days. The source is present in all our separate integrations, making us confident that it is real. A hint of a fading trend appears to be confirmed by observations on Apr. 7.2, when the (preliminary) 850-μm flux density was 2 ± 0.8 mJy. The source was not detected at 450 μm, with a preliminary rms of 10.0 mJy averaged over the first two days. Assuming these fluxes are due to the burst counterpart, they should represent ‘clean’ measures of its intensity, unaffected by scintillation and extinction.”

NOVA SAGITTARII 1998 Visual magnitude estimates: Apr. 1.73 UT, 9.4 (F. Farrell, Christies

Beach, S. Australia); 4.199, 9.0 (A. Pereira, Cabo da Roca, Portugal); 5.165, 9.2 (P. Schmeer, Bischmisheim, Germany); 6.203, 9.7 (Pereira).

1998 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green